OpenStudy (anonymous):

Hi there! Need your urgent help to solve below word problem:- The total amount of pocket money of 4 male students is equal to 3/5 of the pocket money of a female student. What is the average amount of the pocket money of 8 male students and 12 female students if each male student has $36? Thank you.

OpenStudy (anonymous):

ok, you need to find the total amount of money the male students have 8x36=288 now that you have that number, you have to compare it to the ratio 288/x=3/5 cross multiply 3x=1440 simplify x=480 now this x value is the total money the girls have divide 480 by 12 and you get 40, meaning each girl has $40

OpenStudy (anonymous):

if youre looking for the average, add up all of the money and divide by the number of people

OpenStudy (anonymous):

dat means the ans is $38.40? But the child's model answer is $201.60 that puzzles me. please explain. thanks!

OpenStudy (anonymous):

how on earth could each person have $200 when, what? now im confused, what is the model answer?

OpenStudy (anonymous):

the model answer for the average amount is $201.60, i suppose this is the average of 20 students?

OpenStudy (anonymous):

oh wait i misread, the money of 4 males is 3/5 the money of one female

OpenStudy (anonymous):

im getting $158.4, im sorry but im stumped too

OpenStudy (anonymous):

i got that too. Thanks 4 ur effort!